OpenStudy (anonymous):

If T(44)=17 then find lim(h->0) (T(44+h)-T(44))/h

OpenStudy (anonymous):

It was not given... all we got was the table: x 35 44 53 62 71 T (x) 25 17 15 32 28

OpenStudy (anonymous):

I originally put 1 figuring \[\lim_{h \rightarrow 0} \frac{ ((17+h)-17) }{ h } = h/h=1\]

OpenStudy (aum):

They are essentially asking you to find the slope of T at x = 44. You can approximate it by assuming a straight line between x = 44 and x = 53. You can also evaluate it on the left side and perhaps take the average of the two.

OpenStudy (anonymous):

ok, thanks!! I think I know what to do from here...

OpenStudy (aum):

You are welcome.